OpenStudy (anonymous):

Will Fan and Medal A circle has a radius of 12 m. What is the approximate area of the circle? Use 3.14 to approximate pi. Write the answer to the nearest hundredth

OpenStudy (♪chibiterasu):

\[A = \pi r^2\]

OpenStudy (butterflydreamer):

Area of a circle if given by the formula : Area = pi * r^2 OR Area = pi * r * r So we know that r = 12 and pi = 3.14 so just plug these values into the formula above and then into a calculator to find your answer :)

OpenStudy (anonymous):

so 3.14*12*12 ?

OpenStudy (♪chibiterasu):

yep.

OpenStudy (butterflydreamer):

exact;y! ^_^

OpenStudy (anonymous):

452.16?

OpenStudy (butterflydreamer):

yes , that's correct

OpenStudy (anonymous):

Ok thank you both!